OpenStudy (anonymous):

x + (–a) is rational because _______________________. A. it is the sum of two rational numbers. B. it is the sum of two irrational numbers. C. it represents a non-terminating, non-repeating decimal. D. its terms cannot be combined.

OpenStudy (solomonzelman):

it depends. It can be any constants x and a that rational to get A, or irrational (like square cube or any other sorts for example) to get B. It can be d if a or x is an imaginary/complex number.

OpenStudy (solomonzelman):

if a or x (but not both are complex it is then D. See? it can be any of the choices, depending on what course you take.

OpenStudy (anonymous):

What do you mean course?

OpenStudy (solomonzelman):

like have you learned about imaginary numbers. and even if you didn't it is still between A and B.

OpenStudy (solomonzelman):

(imaginary number, \(\normalsize\color{blue}{ \rm i }\) which is equivalent to \(\normalsize\color{blue}{ \rm \sqrt{-1} }\) )

OpenStudy (anonymous):

I learned about it but I still don't get it. I'm doing some summer school online thing and I'm on complex numbers.

OpenStudy (solomonzelman):

As I see the problem, the only thing I see is what you wrote in the question box. Maybe there is more context to the problem than what I see ?

OpenStudy (anonymous):

Do you want me to show you the whole question?

OpenStudy (solomonzelman):

It would be good :)

OpenStudy (anonymous):

Let a be a rational number and b be an irrational number. Assume that a + b = x and that x is rational. Then b = x – a = x + (–a). x + (–a) is rational because _______________________. However, it was stated that b is an irrational number. This is a contradiction. Therefore, the assumption that x is rational in the equation a + b = x must be incorrect, and x should be an irrational number. In conclusion, the sum of a rational number and an irrational number is irrational.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \rm a=rational }\) \(\normalsize\color{blue}{ \rm b=irrational }\) \(\normalsize\color{blue}{ \rm a+b~~=~~x~~=~~rational }\)

OpenStudy (solomonzelman):

I think D.

OpenStudy (anonymous):

Then I'm going to put D and see if that's right.